In January 2007, the average price of an asset was $27,958. 9 years earlier, the average price was $21,808. What was the annual increase in selling price?

3.08%

2.80%

-2.72%

2.52%

3.36%
Answer #1

B. 2.80%

To answer this question, we can use either the FV or the PV formula. Both will give the same answer since they are the inverse of each other. We will use the FV formula, that is:

FV = PV(1 + r)t

Solving for r, we get:

r = (FV / PV)1 / t – 1

r = ($27,958 / $21,808)1/9 – 1 = 2.80%